Fermi National Laboratory
 

Accelerator Update for Jun 16, 2006 - June 19, 2006

 

Friday June 16

The day shift began with Operations monitoring store 4772 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Booster, Main Injector (MI), and Recycler conducting studies, and with NuMI and MiniBooNE taking beam.

Switchyard began an hour study at 1:20 PM.

Operations began preparing the Recycler at 2:48 PM to stash antiprotons.

TeV experts began an end-of-store study at 3:19 PM.

The stash ended and Pbar resumed stacking at 4:35 PM.

Operations terminated store 4772 at 5:06 PM.

Operations prepared D0 for an access at 5:23 PM.

Operations put the Tevatron (TeV) into a dry squeeze at 6:43 PM.

Operations put the TeV into shot setup at 7:14 PM.

Operations established store 4774 at 9:19 PM with an initial luminosity of 95.13E30. Pbar experts began a study period.

Pbar resumed stacking at 10:42 PM.

Saturday June 17

The midnight shift began with Operations monitoring store 4774 and the stash, with Pbar stacking, with CDF and D0 using the colliding beam, and with MiniBooNE and NuMI taking beam.

At 1:22 AM, store 4774 aborted. There was no quench. Operators investigated and discovered a problem with the TeV sector B2 dump switch.

Pbar experts completed their studies at 3:37 AM.

The Recycler began preparing to stash 4:38 AM, but then put the stash on hold until after the TeV was put into a dry squeeze.

Operations put the TeV into a dry squeeze at 5:23 AM. Recycler resumed preparing to stash.

The stash ended and Pbar resumed stacking at 6:54 AM.

At 8:07 AM, a Linac RF station (KRF4) tripped off. Operators managed to reset the station at 8:29 AM.

The Recycler began preparing to stash at 11:02 AM. The stash ended and Pbar resumed stacking at 12:04 PM.

At 12:41 PM, Booster began having beam position problems. Operations called an expert who thought there was a tuning problem in the 400MeV area. Operations called in an expert.

Operations put the TeV into shot setup at 1:47 PM.

Operations established store 4778 at 5:35 PM with an initial luminosity of 55.35E30. Pbar resumed stacking.

The Recycler began preparing to stash at 9:20 PM.

At 9:33 PM, the 400 MeV Debuncher water skid tripped off. It reset okay.

Operations halted the stash at 10:25 PM after three transfers due to the poor transfer efficiency.

At 11:05 PM, the 400 MeV Debuncher water skid tripped off for the first of four times in thirty minutes.

At 11:25 PM, Operations reported that they found a Booster RF station (BRF5) off and bypassed. Operators could find no reason for this, so they returned it to service.

At 11:58 PM, the 400 MeV Debuncher water skid tripped off for the first of two times in an hour.

Sunday June 18

The midnight shift began with Operations monitoring store 4778 and the stash, with Pbar stacking, with CDF and D0 using the colliding beam, and with MiniBooNE and NuMI taking beam.

The Recycler began preparing to stash at 2:38 AM.

At 2:41 AM, Operation reported that the 400 MeV Debuncher water skid had tripped off four times in fifteen minutes. It reset each time.

The stash ended and Pbar resumed stacking at 3:35 AM.

At 3:43 AM, the 400 MeV Debuncher water skid tripped off. It reset okay.

The Recycler began preparing to stash at 9:08 AM. The stash ended and Pbar resumed stacking at 9:54 AM.

The Recycler began preparing to stash at 2:49 PM. The stash ended and Pbar resumed stacking at 3:55 PM.

Operations terminated store 4778 at 4:09 PM.

Operation put the TeV into shot setup at 4:50 PM.

At 6:13 PM, Operations halted the shot setup. D0 reported that they had a calorimeter power supply fail and needed an access to replace the supply. Operators prepared D0 for access.

The D0 access ended at 7:28 PM.

At 7:53 PM, Operations reported a feed down sextupole power supply trip (C:S3A3A) due to a quench indication. Operations had the supply replaced by 9:04 PM.

Operations put the TeV into shot setup at 9:46 PM.

At 9:52 PM, Operations reported that the MiniBooNE Horn power supply tripped on a ground fault indication. Operators called in an expert. Later the expert reported that the repairs would be made on Monday day shift.

At 11 PM, a Pbar power supply (M:HV102) tripped off due to a LCW leak and wouldn’t reset. This also caused the loss of the Pbar beam permit, which in turn held off NuMI. Operations called in an expert.

At 11:42 PM, Operations changed the Time Line Generator (TLG) so that NuMI could resume taking beam.

Monday June 19

The midnight shift began with the TeV in shot setup, with Pbar off due to HV102, with CDF and D0 waiting for the colliding beam, with MiniBooNE off due to a Horn power supply problem, and with NuMI taking beam.

Operations established store 4781 at 12:35 AM with an initial luminosity of 103.9E30.

A power supply expert reported at 1:15 AM that he had replaced a LCW hose on HV102 and put a fan on the supply so it would dry out.

The Recycle began preparing to stash around 1:30 AM.

At 1:41 AM, Operations reported that M:HV202 was off and would not turn on; they investigated and found that the supply was tied, by interlocks, to HV102.

The stash ended at 2:29 AM. HV102 still wouldn’t reset so there was no stacking.

The Plans for Monday

The plans for today are to stack, stash, and store. D0 has a problem that will need about a 16-hour access to repair. Tentatively, the Run Coordinator has planned this to begin at 6:30 AM on Tuesday, 6/20/06. TeV accesses may also be allowed.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/19/2006    email Fermilab